OpenStudy (anonymous):

In a quadrilateral MNOP the length of side MN is (3x - 7) units and the length of side ON is (x + 13) units. What value of x will make it a rhombus? A) 4 B) 5 C) 8 D) 10

OpenStudy (mertsj):

You know that all sides of a rhombus are equal, right?

OpenStudy (anonymous):

Rhombus, all sides are equal, set 3x-7=x+13

OpenStudy (cornzyblack):

Since a property of a rhombus is that the 4 sides are of equal length then, we can say that MN = ON therefore, 3x - 7 = x + 13 taking like terms then, 3x - x = 13 + 7 2x = 20 dividing both sides by 2 then , x = 10 so x must be 10